In 2007​,Amazon reported total revenue of ​$93.6 billion. Ten years later it generated ​$483.2billion in revenue. What was the compound growth rate in revenue from 2007 to 2017​?

The annual compound growth rate of net income is_____​%. ​(Round to two decimal​ places.)
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

483.2=93.6*(1+r/100)^10

(483.2/93.6)^(1/10)=(1+r/100)

(1+r/100)=1.1784

r=1.1784-1

=17.84%(Approx).